# Enhanced Generative Data Augmentation for Semantic Segmentation via Stronger Guidance

### v2

@@ Abstract
Data augmentation is a widely used technique for creating training data for tasks that require labeled data, such as semantic segmentation. This method benefits pixel-wise annotation tasks requiring much effort and intensive labor. Traditional data augmentation methods involve simple transformations like rotations and flips to create new images from existing ones. However, these new images may lack diversity along the main semantic axes in the data and not change high-level semantic properties. To address this issue, generative models have emerged as an effective solution for augmenting data by generating synthetic images. Controllable generative models offer a way to augment data for semantic segmentation tasks using a prompt and visual reference from the original image. However, using these models directly presents challenges, such as creating an effective prompt and visual reference to generate a synthetic image that accurately reflects the content and structure of the original. In this work, we introduce an effective data augmentation method for semantic segmentation using the Controllable Diffusion Model. Our proposed method includes efficient prompt generation using Class-Prompt Appending and Visual Prior Combination to enhance attention to labeled classes in real images. These techniques allow us to generate images that accurately depict segmented classes in the real image. In addition, we employ the class balancing algorithm to ensure efficiency when merging the synthetic and original images to generate balanced data for the training dataset. We evaluated our method on the PASCAL VOC datasets and found it highly effective for synthesizing images in semantic segmentation.
@@ 1 INTRODUCTION
Figure 1: The controllable data augmentation pipeline for the semantic segmentation task combines our proposed methods: Class-Prompt Appending and Visual Prior Combination
@@
An alternative to enhancing data diversity without annotating a new dataset is data augmentation, which creates more training examples by leveraging an existing dataset. Commonly used data augmentation methods in semantic segmentation include rotating, scaling, flipping, and other manipulations of individual images. These techniques encourage the model to learn more invariant features, thereby improving the robustness of the trained model. However, basic transformations do not produce novel structural elements, textures, or changes in perspective. Consequently, more advanced data augmentation methods utilize generative models for different tasks, such as those described in the following works [Trabucco et al., 2024, Azizi et al., 2023, He et al., 2023, Fang et al., 2024, Wu et al., 2024]. Generative models leverage the ability to create new images based on inputs such as text, semantic maps, and image guidance (edge, line art, soft edge) to specific tasks and data augmentation needs. Notably, Stable Diffusion (SD) [Rombach et al., 2022, Podell et al., 2024] and its variants [Zhang et al., 2023, Mou et al., 2023] excel in producing high-quality images based on input conditions.
@@
The segmentation mask annotation can be easily computed in data augmentation using simple transformations. However, with generative models, data augmentation is more problematic as it requires generating new images while still matching the segmentation mask. A simple approach to the aforementioned challenge is to use Inpainting model [Rombach et al., 2022] to change the labeled classes in the images while preserving the remaining information. Although it can increase the diversity of the data classes, the diversity of the surrounding elements is limited when applying the method above. In addition, some studies [Mou et al., 2023, Zhang et al., 2023, Chae et al., 2023] propose image generation models that can be controlled from the segmentation mask of the image. However, all the above-mentioned methods require training on semantic segmentation datasets for generative models, which results in the model being unable to generate classes that are not part of the trained dataset.
@@
With a deep understanding of pre-trained generative models, we can identify their strengths with a vast knowledge base, such as generality and good controllability. In this paper, we propose using controllable generative models with prior visual without training on semantic segmentation datasets to augment data for semantic segmentation. The proposed methods ensure that the generated images match the original images in terms of the number of labels and their localizations but with transformations in color, context, and style. To achieve high performance in data augmentation, we propose Visual Prior Combination and Class-Prompt Appending to enhance the visual representation of labeled classes. This combination method is depicted in Figure 1. In addition, we also use class balancing algorithm to control the number of classes during image generation so that when combining the synthetic dataset and the original data, the classes in the extended dataset are more balanced.
@@ 2 RELATED WORK
Figure 2: Common issues of the generated prompt. Red describes the missing classes in the generated prompt while blue marks the ones appearing in the sentence.
@@
Data augmentation using generative models is commonly applied in classification tasks [Trabucco et al., 2024, Azizi et al., 2023, He et al., 2023]. Unlike classification tasks, to augment data for Segmentation or Object Detection tasks, where the synthetic images must ensure the location of the objects, Inpainting model [Rombach et al., 2022] is considered an option because it allows to specify what to edit as a mask. However, because it is only possible to edit objects in the mask, the contexts around the object are almost kept the same, which leads to the synthetic images needing to be more diverse in terms of content.
@@
Some studies [Mou et al., 2023, Zhang et al., 2023, Chae et al., 2023] suggest using semantic segmentation maps to guide image generation. This means labeling each pixel to show which class it belongs to, helping to create accurate images with correct object locations and details. However, these methods need to be trained on segmentation datasets. This limits the synthetic images to only the trained classes and makes them unclear for untrained classes. For example, the method in [Mou et al., 2023, Zhang et al., 2023] is trained on the ADE20K [Zhou et al., 2019] dataset, which has various images from different contexts like indoor, outdoor, industrial, and natural scenes. Suppose we need to generate synthetic images for the BDD100K dataset [Yu et al., 2020], which includes images from vehicle cameras. In that case, the model can not generate new classes not in the ADE20K dataset, such as traffic signs, traffic lights, and road markings. Because of this limitation, we do not use models guided by semantic segmentation maps in this study to keep our method general.
@@
Instead of directly selecting synthetic images to train the model, some studies [Fang et al., 2024, Wu et al., 2024] propose using post-filtering techniques to choose the best synthetic images. However, selecting one high-quality image from many can be time-consuming, and imperfect filtering can lead to a low-quality synthetic dataset. In addition, with the ability to generate different images from the same input and change the random seeds, selecting the best ones based on multiple results does not accurately reflect the image generation ability. Therefore, in this paper, we directly use the generated images without going through any post-filtering techniques to demonstrate the effectiveness of the proposed method. In addition, in Section 4.3, we also integrate the filter using CLIP Encoder [Fang et al., 2024] to demonstrate the compatibility of the proposed method when applying filters.
@@
(b) The results of the Visual Prior Combination method demonstrate that previously edge-deficient objects are now fully detailed, and the edge features of the labeled classes are more prominent compared to the background, thereby enhancing the generative model’s attention capability.
Figure 3: The edge feature results before (a) and after (b) using the Visual Prior Combination method..
@@ 3 METHOD
Our proposed image data augmentation pipeline is shown in Figure 1, which consists of three main components: (1) Text prompt construction, (2) Visual Prior Combination, and (3) Controllable Diffusion Generation. Class-Prompt Appending append “class prompt” including the classes visible in the image with “caption” generated from the Image Caption model. Visual Pre-Matching combines the visual pre-information of the real image and the segmentation map. The results of the above two methods are fed into the Controllable Steady Diffusion model to generate the synthetic image. In addition, to generate synthetic data from a given dataset, we using class balancing algorithm to generate data with even distribution among classes. Next, we provide a detailed description of each component.
@@ 3.1 Preliminary
The diffusion models comprise both forward and reverse processes [Ho et al., 2020]. In the forward process, a Markovian chain is defined with noise added to the clean image $x_{0}$:
@@ 3.2.1 Text prompt construction
To generate an image containing the labeled classes as in the original image, we need a robust prompt that describes the original image well to serve as input to the SD model. One way to do this is by constructing a prompt based on the labeled classes. For example, if image $\mathcal{I}_{i}$ contains target classes $\mathcal{C}_{i}$ = [$c_{1}$,…, $c_{M}$], where $M$ is the number of classes in image $\mathcal{I}_{i}$, we can construct a simple prompt such as: “$c_{1}$,…, $c_{M}$” or “An image of $c_{1}$,…, $c_{M}$”. However, using a prompt that solely lists target classes doesn’t clearly describe the original image’s layout. In this case, we can employ the existing or generated captions from the training images in these datasets as text prompts for Stable Diffusion. For example, we can use the provided captions when using the COCO dataset [Lin et al., 2014]. However, most datasets, such as PASCAL VOC [Everingham et al., 2015], BDD100K [Yu et al., 2020], and ADE20K [Zhou et al., 2019], do not have captions, while annotating captions for images also requires intensive labor. Therefore, we propose using an image captioning model such as BLIP-2 [Li et al., 2023] to generate captions for each image. However, we encountered some issues with the generated annotations; many did not include all the actual classes in the images (as illustrated in Figure 2). This issue leads to missing class names in the captions.
@@
To address the abovementioned limitations, we propose combining the generated captions with the class labels of the images, as suggested in [Nguyen et al., 2023]. With an image $\mathcal{I}_{i}$, we append the generated captions $\mathcal{P}_{i}^{g}$ with the class labels $\mathcal{P}_{i}^{c}$ to generate new text prompts $P^{*}_{i}$. This process, known as class-prompt appending, can be represented as: $\mathcal{P}^{*}_{i}$ = “$\mathcal{P}_{i}^{g}$; $\mathcal{P}_{i}^{c}$”. For example, in the sub-figure in the top-right corner of Figure 2, the prompt generated by our proposed method would be “a pink plane on the tarmac; aeroplane, person”. Our method ensures that new text prompts include both general information and the target classes of the original image. This technique helps the synthetic image to have a clear layout similar to the original one and also addresses the problem of missing labeled classes in the synthetic image.
3.2.2 Visual Prior Combination
Unlike Stable Diffusion models that typically only use text prompts to generate images, Controllable Generation models require additional input guidance (Canny Edge [Canny, 1986], Sketch-Guided [Su et al., 2021], Line-Art Edge [Chan et al., 2022], Depth Map [Ranftl et al., 2022], HED soft edge [Xie and Tu, 2015]) generated from the visual prior to determine the image layout. Canny, line-art edge and sketch are the visual priors we propose using to balance the diversity of the image and the precise localization of the generated classes in the image. Our default visual prior is Line-Art Edge to generate visual guidance. In Section 4.3.5, we also discuss other types of visual priors to see how effectively they augment the data. Using the T2I-Adapter [Mou et al., 2023] model, the Linear Detector converts the image into visual prior in-line drawings for each input image. Then, the Adapter generates different resolution features, performing conditional operations at each time step with the UNet denoiser’s features.
In general, methods such as Line-Art Edge, Canny Edge, HED soft edge all suffer from the limitation that the labeled classes in the image may be blurred or small in size, leading to inaccuracies in describing the shape of the labeled classes within the conditional image. Figure 3(a) shows an image produced using Line Art Detection. However, the edge results in this case is missing some details of the person, and the TV monitor is almost absent. The red shapes indicate the missing details in the image. This loss of information also occurs when using HED or Canny edge. These weaknesses result in mislabeling in the synthetic image compared to the original image. We observe that the segmentation labels divide the regions of the labeled so their prior visualization accurately describes the label information in the original image. Therefore, the combination of the prior visual image $\mathcal{I}_{i}$ ($\mathcal{V}^{\mathcal{I}}_{i}$) and prior visual segmentation label $\mathcal{S}_{i}$ ($\mathcal{V}^{\mathcal{S}}_{i}$) ensures that the generated image has a clear layout and well preserves the labeled information. Our proposed combines $\mathcal{V}^{\mathcal{I}}_{i}$ and $\mathcal{V}^{\mathcal{S}}_{i}$ by a weighted sum:
With $\omega_{1}$, $\omega_{2}$ being the trade-off scales when combining $\mathcal{V}^{\mathcal{I}}_{i}$ and $\mathcal{V}^{\mathcal{S}}_{i}$. This combination results in a prior visual that is clear in content and complete information about labeled classes. Figure 3(b) shows an example of how the Visual Prior Combination method can be used to highlight classes in a dataset.
3.3 Create class balancing dataset
@@ 15                  $\mathcal{D}_{gen}\leftarrow\mathcal{D}_{gen}\cup\{\mathcal{I}_{gen}\}$
Table 1: Comparison in mIoU on val set between training of models on real training set and extended training set
@@
To address the issue of class imbalance during model training, we aim for the final dataset $D_{final}$, which merges the original dataset $D_{ori}$ and the synthetic dataset $D_{gen}$, to have a balanced distribution among classes. To create a balanced dataset from the original dataset $D_{ori}$, we use the class balancing algorithm to generate the dataset $D_{gen}$ based on the balancing factor $n_{balance}$, as presented in Algorithm 1. The algorithm consists of three main stages: Initialization, Sorting, and Balancing. In Stage 1, a dictionary $\mathcal{M}$ is initialized to map each class to its associated images. Each image $\mathcal{I}$ is linked to a list of classes it contains. In the next stage, images are arranged in ascending order based on the number of classes they represent, prioritizing those with fewer classes to be generated to maintain the balance. In the final stage, additional images are generated for each class until they reach $n_{gen}$, ensuring an even distribution across classes. This process ensures the dataset is balanced, preventing overrepresentation of any class and promoting more robust model training.

### v3

@@ Abstract
Data augmentation is crucial for pixel-wise annotation tasks like semantic segmentation, where labeling requires significant effort and intensive labor. Traditional methods, involving simple transformations such as rotations and flips, create new images but often lack diversity along key semantic dimensions and fail to alter high-level semantic properties. To address this issue, generative models have emerged as an effective solution for augmenting data by generating synthetic images. Controllable Generative models offer data augmentation methods for semantic segmentation tasks by using prompts and visual references from the original image. However, these models face challenges in generating synthetic images that accurately reflect the content and structure of the original image due to difficulties in creating effective prompts and visual references. In this work, we introduce an effective data augmentation pipeline for semantic segmentation using Controllable Diffusion model. Our proposed method includes efficient prompt generation using Class-Prompt Appending and Visual Prior Blending to enhance attention to labeled classes in real images, allowing the pipeline to generate a precise number of augmented images while preserving the structure of segmentation-labeled classes. In addition, we implement a class balancing algorithm to ensure a balanced training dataset when merging the synthetic and original images. Evaluation on PASCAL VOC datasets, our pipeline demonstrates its effectiveness in generating high-quality synthetic images for semantic segmentation. Our code is available at this https URL.
@@ 1 INTRODUCTION
Figure 1: The controllable data augmentation pipeline for the semantic segmentation task combines our proposed methods: Class-Prompt Appending and Visual Prior Blending
@@
An alternative to enhancing data diversity without annotating a new dataset is data augmentation, which creates more training examples by leveraging an existing dataset. Commonly used data augmentation methods in semantic segmentation include rotating, scaling, flipping, and other manipulations of individual images. These techniques encourage the model to learn more invariant features, thereby improving the robustness of the trained model. However, basic transformations do not produce novel structural elements, textures, or changes in perspective. Consequently, more advanced data augmentation methods utilize generative models for different tasks [Trabucco et al., 2024, Azizi et al., 2023, He et al., 2023, Fang et al., 2024, Wu et al., 2024]. Generative models leverage the ability to create new images based on inputs such as text, semantic maps, and image guidance to specific tasks and data augmentation needs. Notably, Stable Diffusion (SD) models [Rombach et al., 2022, Podell et al., 2024] propose a method for conditional image generation that fusions textual information, bounding boxes, or segmentation masks to generate or inpaint images. Controllable Models [Zhang et al., 2023, Mou et al., 2023] further enhance the guided image generation capability by utilizing visual priors such as edges, depth, segmentation mask, human poses, etc.
@@
The segmentation mask annotation can be easily computed in data augmentation using simple transformations (e.g. translation, scaling, flip). However, with generative models, data augmentation is more problematic as it requires generating new images while still matching the ground truth annotations. A straightforward approach to this challenge is to utilize the Inpainting model [Rombach et al., 2022] to change the labeled regions in the images while keeping the rest of the remaining information. Although this method can enhance the diversity of labeled data, it does not ensure that the newly generated object matches the original structure, and the surrounding elements may lack diversity since they are left unchanged. Additionally, [Mou et al., 2023, Zhang et al., 2023, Chae et al., 2023] propose image generation models that can be controlled via segmentation masks, making these methods highly suitable for efficient data augmentation in semantic segmentation tasks. However, generative models require training on semantic segmentation datasets, which limits their capacity to provide information beyond the scope of the training dataset.
@@
To address the issue of data augmentation for semantic segmentation using generative models, we can leverage a deep understanding of generative models to identify their strengths, such as their broad knowledge base, generalization capabilities, and structural control. In this paper, we propose using Controllable Generative models with prior visual without training on semantic segmentation datasets to augment data for semantic segmentation. However, utilizing Controllable Generative models directly may present challenges, such as generated objects not strictly adhering to their original structure or a lack of labeled classes in the generated images. So, how can we overcome these limitations? In this work, we propose Visual Prior Blending to enhance the visual representation of labeled classes and Class-Prompt Appending to construct text prompts that generate images containing all labeled classes. The combination of these two proposed methods, as shown in Figure 1, enables more effective data augmentation when utilizing Controllable Generative models, resulting in improved performance.
In this work, we analyze and propose methods to address these challenges when generating images using controllable generative models.
Our main contributions are:
  • Class-Prompt Appending: By combining “generated caption” and “labeled classes”, we use this method to construct text prompts containing comprehensive information about the image and its classes. This ensures that all labeled classes are fully represented in the synthesized image.
  • Visual Prior Blending: This method combines the visual priors from the original image and the segmentation masks. It aims to balance the information between the original image and the labeled objects, ensuring that the generated objects maintain the structure defined by the segmentation labels.
  • In addition, we also use a class balancing algorithm to control the number of classes during image generation so that when combining the synthetic dataset and the original data, the classes in the extended dataset are more balanced.
  • We evaluate our proposed method on the VOC7/VOC12 datasets with various settings. The results demonstrate the effectiveness of the proposed approach, particularly in scenarios with limited data samples.
@@ 2 RELATED WORK
Data augmentation using generative models is commonly applied in classification tasks [Trabucco et al., 2024, Azizi et al., 2023, He et al., 2023]. Unlike classification tasks, to augment data for Segmentation or Object Detection tasks, where the synthetic images must ensure the location of the objects. Inpainting model [Rombach et al., 2022, Podell et al., 2024, Yang et al., 2023] is considered an option because it allows to specify what to edit as a mask. However, using an off-the-shelf inpainting model for data augmentation in semantic segmentation tasks may lead to shortcomings. There is no guarantee that the newly created objects align with the ground truth annotations, and the model may also tend to replace smaller objects with background elements.
Figure 2: Classes missing in the generated prompts: Red describes the missing classes in the generated prompt while blue marks the ones appearing in the sentence.
Figure 3: Common issues of using generated captions and simple text prompts: In addition to generated prompts and simple text prompts, the four visualizations include: original images with labeled classes, guidance images (line art), images generated by simple text prompts, and images generated by generated prompts.
@@
Early studies in synthetic data generation [Mou et al., 2023, Zhang et al., 2023, Chae et al., 2023] propose using semantic segmentation maps to guide image generation and create an efficient solution for data augmentation for semantic segmentation. Labeling each pixel to show which class it belongs to helps to create accurate images with correct object locations and details. However, these methods require training on specific segmentation datasets, which limits the generation of synthetic images for classes not included in the training data. For example, [Mou et al., 2023, Zhang et al., 2023] propose models trained on the ADE20K dataset [Zhou et al., 2019], which has various images from different contexts like indoor, outdoor, industrial, and natural scenes. When generating synthetic images based on segmentation masks from the BDD100K dataset [Yu et al., 2020], which consists of images captured from car dashcams in self-driving scenarios, the model is unable to generate classes not present in the ADE20K dataset. These missing classes include traffic signs, traffic lights, and lane markings. Due to this limitation, we avoid using models guided by semantic segmentation maps in this study to maintain the generality of our method.
@@
Instead of directly selecting synthetic images to train the model, some studies [Fang et al., 2024, Wu et al., 2024] propose using post-filtering techniques to choose the best synthetic images. However, selecting one high-quality image from many can be time-consuming, and imperfect filtering can lead to a low-quality synthetic dataset. In addition, with the ability to generate different images from the same input and change the random seeds, selecting the best ones based on multiple results does not accurately reflect the image generation ability. Therefore, in this paper, we directly use the generated images without going through any post-filtering techniques to demonstrate the effectiveness of the proposed method. In addition, in Section 4.3.1, we also integrate the filter using CLIP Encoder [Fang et al., 2024] to demonstrate the compatibility of the proposed method when applying filters.
@@
(b) The results of the Visual Prior Blending method demonstrate that previously edge-deficient objects are now fully detailed, and the edge features of the labeled classes are more prominent compared to the background, thereby enhancing the generative model’s attention capability.
Figure 4: The edge feature results before (a) and after (b) using the Visual Prior Blending method..
@@ 3 METHOD
Our proposed image data augmentation pipeline is shown in Figure 1, which consists of three main components: (1) Text prompt construction, (2) Visual Prior Blending, and (3) Controllable Diffusion Generation. Class-Prompt Appending append “class prompt” including the classes visible in the image with “caption” generated from the Image Captioning model. Visual Prior Blending method combines the visual pre-information of the real image and the segmentation map. The results of the above two methods are fed into the Controllable Steady Diffusion model to generate the synthetic image. In addition, to generate synthetic data from a given dataset, we use class balancing algorithm to generate data with even distribution among classes. Next, we provide a detailed description of each component.
@@ 3.1 Preliminary
Diffusion models comprise both forward and reverse processes [Ho et al., 2020]. In the forward process, a Markovian chain is defined with noise added to the clean image $x_{0}$:
@@ 3.2.1 Text prompt construction
To generate an image containing the labeled classes as in the original image, we need a robust prompt that describes the original image well to serve as input to the SD model. A simple way to do this is by constructing a prompt based on the labeled classes. For example, if image $\mathcal{I}_{i}$ contains target classes $\mathcal{C}_{i}$ = [$c_{1}$,…, $c_{M}$], where $M$ is the number of classes in image $\mathcal{I}_{i}$, we can construct a simple prompt such as: “$c_{1}$,…, $c_{M}$” or “A photo of $c_{1}$,…, $c_{M}$”. However, using a prompt that only lists target classes may not clearly describe the original image’s layout. To improve this, we can use the existing or generated captions from the training images in these datasets as text prompts for SD. For example, we can use the provided captions when using the COCO dataset [Lin et al., 2014]. However, most datasets, such as PASCAL VOC [Everingham et al., 2015], BDD100K [Yu et al., 2020], and ADE20K [Zhou et al., 2019], do not have captions, while annotating captions for images also requires intensive labor. Therefore, we propose using an Image Captioning model such as BLIP-2 [Li et al., 2023] to generate captions for each image. However, image captions have limitations compared to simple prompts, as they often omit some of the actual classes present in the image (as shown in Figure 2). This issue results in missing class names in the captions when using generated descriptions.
@@
Illustrated in Figure 3, some examples demonstrate using simple text prompts and generated captions to create synthetic images through the Controllable SD model. The results indicate that using simple text prompts produces images with messy layouts that do not match the original while using generated prompts leads to missing classes in the generated images due to their absence in the generated captions. However, we observe that these two methods can complement each other’s weaknesses. Therefore, we propose combining generated captions with the image’s class labels to address these limitations. With an image $\mathcal{I}_{i}$, we append the generated captions $\mathcal{P}_{i}^{g}$ with the class labels $\mathcal{P}_{i}^{c}$ to generate new text prompts $P^{*}_{i}$. This process, known as Class-Prompt Appending [Nguyen et al., 2023], can be represented as: $\mathcal{P}^{*}_{i}$ = “$\mathcal{P}_{i}^{g}$; $\mathcal{P}_{i}^{c}$”. For example, in the sub-figure in the top-right corner of Figure 2, the prompt generated by our proposed method would be “a pink plane on the tarmac; aeroplane, person”. Our method ensures that new text prompts include both general information and the target classes of the original image. This technique helps the synthetic image to have a clear layout similar to the original one and also addresses the problem of missing labeled classes in the synthetic image.
@@ 15                  $\mathcal{D}_{gen}\leftarrow\mathcal{D}_{gen}\cup\{\mathcal{I}_{gen}\}$
3.2.2 Visual Prior Blending
Table 1: Semantic Segmentation Evaluation: Comparison in mIoU (%) on val set between models’ training on the original training set ($D_{origin})$ and the extended training set ($D_{final}$).
@@ 3.2.2 Visual Prior Blending
Unlike Stable Diffusion models that typically only use text prompts to generate images, Controllable Generation models require additional input guidance (Canny Edge [Canny, 1986], Sketch-Guided [Su et al., 2021], Line-Art Edge [Chan et al., 2022], Depth Map [Ranftl et al., 2022], HED soft edge [Xie and Tu, 2015]) generated from the visual prior detector to determine the image layout. Canny, Line-Art Edge, and Sketch are the visual priors we propose using to balance the diversity of the image and the precise structure of the generated classes in the image. Our default visual prior is Line-Art Edge to generate visual guidance. In Section 4.3.5, we also discuss other types of visual priors to see how effectively they augment the data. Using the T2I-Adapter [Mou et al., 2023] model, the Line Art Detector converts the image into visual prior in-line drawings for each input image. Then, the Adapter generates different resolution features, performing conditional operations at each time step with the UNet denoiser’s features.
In general, methods such as Line-Art Edge, Canny Edge, or HED soft edge all suffer from the limitation that the labeled classes in the image may be blurred or small in size, leading to inaccuracies in describing the structure of the labeled classes within the conditional image. Figure 4(a) shows an image produced using Line Art Detection. However, the edge results in this case are missing some details of the person, and the TV monitor is almost absent. The red shapes indicate the missing details in the image. This loss of information also occurs when using HED or Canny Edge. These weaknesses result in mislabeling in the synthetic image compared to the original image. We observed that although the segmentation labels of real images cannot fully describe an image’s content, they provide accurate information about the labeled classes. Based on this observation, we propose combining the real image’s prior visualization with the labels before feeding them into the controlled image generation model. The combination of the prior visual image $\mathcal{I}_{i}$ ($\mathcal{V}^{\mathcal{I}}_{i}$) and prior visual segmentation label $\mathcal{S}_{i}$ ($\mathcal{V}^{\mathcal{S}}_{i}$) ensures that the generated image has a clear layout and well preserves structures the class labeled information. Our proposed combines $\mathcal{V}^{\mathcal{I}}_{i}$ and $\mathcal{V}^{\mathcal{S}}_{i}$ by a weighted sum:
With $\omega_{1}$, $\omega_{2}$ being the trade-off scales when combining $\mathcal{V}^{\mathcal{I}}_{i}$ and $\mathcal{V}^{\mathcal{S}}_{i}$. This combination results in a prior visual that is clear in content and complete information about labeled classes. Figure 4(b) shows how the Visual Prior Blending method can preserve the structure labeled classes in an image.
3.3 Create class-balancing dataset
To address the issue of class imbalance during model training, we aim for the final dataset $D_{final}$, which merges the original dataset $D_{origin}$ and the synthetic dataset $D_{gen}$, to have a balanced distribution among classes. To create a balanced dataset from the original dataset $D_{origin}$, we use the class balancing algorithm to generate the dataset $D_{gen}$ based on the balancing factor $n_{balance}$, as presented in Algorithm 1. The algorithm consists of three main stages: Initialization, Sorting, and Balancing. In Stage 1, a dictionary $\mathcal{M}$ is initialized to map each class to its associated images. Each image $\mathcal{I}$ is linked to a list of classes it contains. In the next stage, images are arranged in ascending order based on the number of classes they represent, prioritizing those with fewer classes to be generated to maintain the balance. In the final stage, additional images are generated for each class until they reach $n_{balance}$, ensuring an even distribution among classes. This process ensures the dataset is balanced, preventing the overrepresentation of any class and promoting more robust model training.